Answer the following questions about layered protocol models. Consider the 5-layer Internet Protocol Stack. The layers...

  1. Answer the following questions about layered protocol models.

    1. Consider the 5-layer Internet Protocol Stack. The layers from top to bottom are application, transport, network, link, and physical. All layers use headers and the transport layer also uses a trailer. Sketch a packet showing headers and trailers.

    2. Assume another protocol stack that uses variable length packets of maximum payload size 250 bytes. In addition to the payload, there are 20 bytes of network-layer header, 20 bytes of transport-layer header, and 10 bytes of trailer in each packet. For a 650-byte message, what is the overhead (in %) for this protocol?
Answer #1

1)

2)

Size of the message = 650 bytes

Maximum payload size = 250 bytes

The 650 byte message can be partitioned as :

Size of 1st  packet = Payload size + Size of network header + Size of transport header + Size of transport trailer

= 250 + 20 + 20 +10

= 300 bytes

Size of 2nd packet = Payload size + Size of network header + Size of transport header + Size of transport trailer

= 250 + 20 + 20 +10

= 300 bytes

Size of 3rd packet = Payload size + Size of network header + Size of transport header + Size of transport trailer

= 150 + 20 + 20 +10

= 200 bytes

So three packets were used to send the 650 byte message

Total data sent including overhead using headers and trailer = 300 + 300 + 200 = 800 bytes

Size of headers and trailer = 150 bytes

Total overhead is